OpenStudy (anonymous):

2x + 24 + y+ z + w = 360 solve for x y z w

OpenStudy (anonymous):

you can solve it using matrix method.

OpenStudy (anonymous):

to solve for x, move all other numbers and variables to the other side. so subtract 24 y z and w from 360 to get 2x=336-y-z-w then to finish solving for x, divide by 2. you can use a similar method to find the individual values of y, w and z

OpenStudy (anonymous):

its for the interior angles of a trapezoid i need to label the measures

OpenStudy (anonymous):

oh, then that is quite different. can you supply a picture?

OpenStudy (anonymous):

no but 2x + 24 is the top right angle bottom right is just x and the whole left side is blank i know that togethor the angles equal 360

OpenStudy (anonymous):

well the base angles of an isoceles trap are congruent, so you can start there. also, opposite angles of a quadrilateral are supplementary, so that should also help you get started :)

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry, okay here we go: 2x+24+x=180 because they are consecutive interior angles (recall that the top and bottom of trapezoids are parallel). solve this, what do you get for x?

OpenStudy (anonymous):

no, you should first combine like terms and you should have 3x=180-24. can you solve it now?

OpenStudy (anonymous):

52

OpenStudy (anonymous):

*x=52

OpenStudy (anonymous):

correct. so x is 52, mening the bottom right angle is 52 and the top right angle is 52+24

OpenStudy (anonymous):

*correction, the top right angle is 52*2+24

OpenStudy (anonymous):

what is the bottom right then

OpenStudy (anonymous):

te bottom right is 52, because x=52

OpenStudy (anonymous):

the top right is 128

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

how do i do the left side it has to equal 180 right

OpenStudy (anonymous):

or does it not matter because it is not labeled

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem :) hope this helped... geometry can be hard, but if you get down those theorems you will be amazed how much everything makes sense :)